OpenStudy (anonymous):

can someone please show me step by step how to solve for x in this problem: 4.5 = 4.74 + log(x/.15)

OpenStudy (dumbcow):

is that natural log or log base 10 ?

OpenStudy (anonymous):

natural log of x divided by .15

OpenStudy (dumbcow):

\[\ln (\frac{x}{.15}) = 4.5 - 4.74 = -.24\] undo the log by taking "e" raised to both sides

OpenStudy (dumbcow):

\[\frac{x}{.15} = e^{-.24}\] \[x = .15 e^{-.24}\]